JOHNSON & JOHNSON AND SUBSIDIARIES CONSOLIDATED STATEMENTS OF EARNINGS (Dollars and Shares in Millions Except Per Share AmounJOHNSON & JOHNSON AND SUBSIDIARIES CONSOLIDATED STATEMENTS OF EQUITY (Dollars in Millions) (Note 1) Accumulated Other Compreh(Dollars In millions) 2018 51 Cost of products sold Selling, marketing and administrative expenses Research and development e1. Summary of Significant Accounting Policies Principles of Consolidation The consolidated financial statements include the apresented in the financial statements. The Companys operating leases will result in the recognition of additional assets andcustomers for retumed goods. The Companys sales retums reserves are accounted for in accordance with the U.S. GAAP guidance2,200 Goodwill as of December 30, 2018 and December 31, 2017, as allocated by segment of business, was as follows: (Dollars iAll derivative instruments are recorded on the balance sheet at fair value. Changes in the fair value of derivatives are recoOther Cost of BAD (Income) The following table is a summary of the activity related to derivatives and hedges for the fiscal7. Borrowings The components of long-term debt are as follows: Effective Rate Effective Rate 31 2017 90] [8 ༡] 1.| 39 $$) 83Question 11 Round your following computation to two decimal places (5.6793 = 5.68). Based on Johnson & Johnsons current ratiQuestion 12 Which of Johnson & Johnsons accounts was increased when the company made a sale on account? a Accounts receivablQuestion 13 When Johnson & Johnson collects an accounts receivable from a customer: na retained earnings increase. Cb accountQuestion 14 Johnson & Johnsons most recent income statement shows the companys results of operations: a. At December 31, 20Question 15 Johnson & Johnsons total net operating revenue for its most recent year end was (in millions): O a. $22,201 ООО
Answer #1

Question 11

Let us begin by calculating the Current ratio for both the years

For 2018 = 46,033 / 31,230 = 1.47 : 1
For 2017 = 43,088 / 30,537 = 1.41 : 1

The ratio has improved marginally

Question 12

The following entry is recorded when a sale is made on account

Debit Accounts Receivable
Credit Revenue

So this means ACCOUNTS RECEIVABLE is increased
When you make a sale on account, the party (the DEBTOR) is obligated to pay. So it increases the receivable balance.

Question 13

Entry would be
Cash Debit
Accounts Receivable Credit

a Retained earning increase See the entry. No account that affects retained earnings is involved
b accounts payable decreases See the entry. No account that affects payable is involved
c total assets increase Cash is increasing the total assets whereas accounts receivable is decreasing the total assets. So in total there is no impact
d Net Income does not change This is the correct answer

Question 14

The correct answer is Option D

An income statement is never as on a particular date. It shows results for a particular period. In this case this period was fiscal period ended on December 30, 2018.

Question 15

The correct answer is Option B. $ 81,581 million. This is the sales to customers figure in the Consolidated Statement of Earnings